OpenStudy (anonymous):

A right triangle is shown. A right triangle with hypotenuse length of 15, base angle of 22 degrees and altitude of x. Which is closest to the value of x ?

OpenStudy (anonymous):

the 15 is on the long side and the 22 is where you put the x

OpenStudy (anonymous):

i dont know how to do that

Vocaloid (vocaloid):

OH WHOOPS HOLD ON, let me draw the diagram again

OpenStudy (anonymous):

ok :)

Vocaloid (vocaloid):



OpenStudy (anonymous):

the 22 is where you put the x

Vocaloid (vocaloid):

nope, "base angle" means the angle on the bottom

Vocaloid (vocaloid):

"altitude" is the height of the triangle

OpenStudy (anonymous):

oh

Vocaloid (vocaloid):

anyway, sin(22) = opposite/hypotenuse = x/15 solve for x

OpenStudy (anonymous):

f 13.9 g 9.0 h 5.6 j 6.1 these are the awncer choices

Vocaloid (vocaloid):

alright, we know that sin(22) = x/15, so what is x?

OpenStudy (anonymous):

um how do i do that ?

Vocaloid (vocaloid):

hint: a = b/c b = c*a

OpenStudy (anonymous):

so do i use a calculator

Vocaloid (vocaloid):

now look at what I wrote before: sin(22) = x/15

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

what do i put into the calculator

Vocaloid (vocaloid):

I'm trying to help you figure it out on your own let's try an easier problem: 2 = x/4, what is x?

OpenStudy (anonymous):

idk i never learned sin and that stuff because i was new to the school and they already tought that

Vocaloid (vocaloid):

let's ignore the sin stuff for now, just solve this problem: 2 = x/4 if x divided by 4 is 2, then what is x?

OpenStudy (anonymous):

2

Vocaloid (vocaloid):

no, please try again

Vocaloid (vocaloid):

what number gives you 2 when you divide it by 4?

OpenStudy (anonymous):

4

Vocaloid (vocaloid):

what is 8 divided by 4?

OpenStudy (anonymous):

2

Vocaloid (vocaloid):

great! so, what number gives you 2 when you divide it by 4?

OpenStudy (anonymous):

8

Vocaloid (vocaloid):

excellent! so, if 2 = x/4, then x = 8, because 8 divided by 4 is 2

Vocaloid (vocaloid):

in order to find x, we multiplied 2*4

Vocaloid (vocaloid):

so, back to the sin stuff if sin(22) = x/15, we can find x by multiplying sin(22) and 15 so what is x?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so would i do something like that to get my awncer to the first question

Vocaloid (vocaloid):

yup! so, if sin(22) = x/15, what is x?

Vocaloid (vocaloid):

what two numbers do you need to multiply to get x?

OpenStudy (anonymous):

15 and 22 ?

Vocaloid (vocaloid):

not quite what is on the left side of the equation?

OpenStudy (anonymous):

22

Vocaloid (vocaloid):

no, the whole thing

OpenStudy (anonymous):

sin(22)

Vocaloid (vocaloid):

great! so, what do you multiply to get x?

OpenStudy (anonymous):

22

Vocaloid (vocaloid):

no, please think again

OpenStudy (anonymous):

sin(22)

Vocaloid (vocaloid):

yes! so, what do you multiply to get x?

OpenStudy (anonymous):

sin(22)

Vocaloid (vocaloid):

you need to multiply TWO things, what's the other number?

Vocaloid (vocaloid):

sin(22), and....?

OpenStudy (anonymous):

15

Vocaloid (vocaloid):

YES!, sin(22) and 15!

Vocaloid (vocaloid):

x = sin(22) times 15

OpenStudy (anonymous):

and that gives me 5.6

Vocaloid (vocaloid):

excellent work, you're done

OpenStudy (anonymous):

thank you so much !!!!!